# Shape-constrained density estimation with Wasserstein projection

### 09 Mar 2026 v1

@@ 1. Introduction
Let $X_{1},\ldots,X_{n}$ be independent samples from an unknown distribution $\mu^{*}$. Let $\mathcal{F}$ be a given set of distributions, possibly infinite dimensional, that encodes the shape constraint; it represents our statistical model. The problem is to come up with an estimator $\hat{\mu}_{n}$ of $\mu^{*}$ with values in $\mathcal{F}$. We allow the model $\mathcal{F}$ to be misspecified, so it is possible that $\mu^{*}\notin\mathcal{F}$. Two important cases that will be studied in detail in this paper are (i) the space of non-decreasing densities on $\mathbb{R}_{+}$ (the monotone case, see Section 3) and (ii) the space of log-concave densities on $\mathbb{R}$ (the log-concave case, see Section 4).
@@
provided that it is well-posed. In the monotone case, this leads to Grenander’s estimator introduced in [15]. The log-concave was studied in [40] if not earlier. Both, and other, shape constraints have been intensively studied in the literature, see [13, 16, 30, 31] and the references therein.
In this paper, we study an alternative approach to the estimation problem, by using optimal transport [32, 38, 39]. For $p\geq 1$, let $\mathcal{W}_{p}(\mu,\nu)$ be the $p$-Wasserstein distance between distributions $\mu$ and $\nu$ on $\mathbb{R}$ (necessary concepts in optimal transport will be reviewed in Section 2.1). Let $\mu_{n}:=\sum_{i=1}^{n}\frac{1}{n}\delta_{X_{i}}$ be the empirical distribution of the data. Our Wasserstein projection estimator is defined by
@@
which may be regarded as a special case of the minimum Kantorovich estimator [5]. In order that the Wasserstein projection estimator exists uniquely (see Theorem 2.8), we assume (i) $p>1$, (ii) distributions in $\mathcal{F}$ have finite $p$-th moment, and (iii) $\mathcal{F}$ is closed with respect to $\mathcal{W}_{p}$ and is displacement convex in the sense of McCann [21]. These conditions not only guarantee that (1.2) is a convex optimization problem, but also accommodate a wide variety of shape constraints. In most of the paper, we restrict further to the case $p=2$, where the Wasserstein projection satisfies a desirable Lipschitz property with respect to $\mathcal{W}_{2}$. From this, finite sample performance of the estimator can be established using known results on convergence of empirical distributions with respect to the Wasserstein distance (Propositions 2.15 and 4.5).
@@
Although these properties appear to be qualitatively similar to those of the maximum likelihood estimators, there are important differences. For example, the support of the estimated density is generally not equal to the convex null of the data points, and the set of break points (of intervals where the density is constant or log-affine) may not be a subset of the data points. To give a quick example (see Example 4.4 for details), suppose that the data $\mu_{n}=\frac{1}{2}\delta_{-1}+\frac{1}{2}\delta_{1}$ is uniformly distributed on $\{-1,1\}$, and $\mathcal{F}$ is the set of log-concave distributions. It is well known that the maximum likelihood estimator returns the uniform distribution $\mathrm{Unif}(-1,1)$. Interestingly, the $2$-Wasserstein projection estimator returns $\mathrm{Unif}(-3/2,3/2)$, which has a wider support. Even in our univariate setting, the proofs of the above structural results are significantly more delicate than those of the classical counterparts. Intuitively, this is because optimal transport incorporates the geometry of the state space. Extending these results to the multivariate setting is interesting but requires different techniques due to the curvature of the Wasserstein space.
@@
Our work contributes to the growing literature of statistical inference based on optimal transport, [^2] This is related to, but different from, inference of transport-based quantities (such as optimal transport maps and Wasserstein barycenters) from samples, and modeling of distributional data. as opposed to classical methods especially maximum likelihood. Recently, an intriguing framework of Wasserstein information geometry has been developed in the parametric setting [19, 3, 2]; of fundamental importance are the Wasserstein information matrix and the resulting Wasserstein–Cramér–Rao inequality [23, 35]. It is interesting to understand to what extend this theory extends to semiparametric and nonparametric settings. Density estimation under Wasserstein distance has been studied by many authors. In particular, minimax estimation of smooth densities in Wasserstein distance was studied in [22, 41], but without shape constraints. Note that smoothness of the density is not assumed in this paper. Statistical properties of the Wasserstein projection estimator have been investigated in one-dimensional location-scale models [1] and circular models [24]. Shape constraints were considered in [11], where the author adopted a computational approach and did not prove structural results of the kind obtained here.
The rest of the paper is organized as follows. In Section 2, we introduce the Wasserstein projection estimator after recalling some basic concepts in optimal transport, and establish general properties including consistency and affine equivariance. In Sections 3 and 4, we specialize respectively to monotone and log-concave density estimation, and prove our main results about structural properties of the Wasserstein projection estimator. Section 5 implements tractable discretizations of Wasserstein projection estimation in the monotone and log-concave settings. In Section 6 we conclude and discuss directions for future research.
@@ 2.1. Wasserstein geometry of univariate distributions
When it is convenient to refer to the boundary values (e.g. when $\mu$ is compactly supported or equivalently $Q_{\mu}$ is bounded), we take $Q_{\mu}(0):=\lim_{u\downarrow 0}Q_{\mu}(u)$ and $Q_{\mu}(1):=\lim_{u\uparrow 1}Q_{\mu}(u)$. By construction, we have
@@
where $(Q_{\mu},Q_{\nu})$ denotes the mapping $u\mapsto(Q_{\mu}(u),Q_{\nu}(u))$ from $(0,1)$ to $\mathbb{R}^{2}$. Then $\pi$ is optimal for $\mathcal{W}_{p}(\mu,\nu)$. In fact, by [32, Theorem 2.9], the comonotonic coupling is the unique optimal coupling when $p>1$.
@@
where $\Pi_{t}(x,y):=(1-t)x+ty$. By construction, the quantile function of $\mu_{t}$ is $Q_{\mu_{t}}=(1-t)Q_{\mu_{0}}+tQ_{\mu_{1}}$. Thus, a set $\mathcal{F}\subset\mathcal{P}(\mathbb{R})$ is displacement convex in the sense of McCann if and only if $\mathcal{Q}_{\mathcal{F}}$ is convex.
@@
It can be shown that $\mu\in\mathcal{F}$ if and only if $\mu$ has a density $f$ which is bounded below and above on its support: $m\leq f(x)\leq M$. It is easy to see that $\mathcal{F}$ is displacement convex.
@@
Let $\mathcal{F}=\{\delta_{m}:m\in\mathbb{R}\}$ be the set of point masses. Clearly, $\mathcal{F}$ is displacement convex. Let $\mu=\frac{1}{2}\delta_{0}+\frac{1}{2}\delta_{1}$. It is easy to verify that
@@
$\begin{split}\mathcal{W}_{p}^{p}(T_{\#}\nu,T_{\#}\mu)=\int_{[0,1]}|Q_{T_{\#}\nu}(u)-Q_{T_{\#}\mu}(u)|^{p}\mathrm{d}u=|b|^{p}\mathcal{W}_{p}^{p}(\nu,\mu).\end{split}$
@@ 2.2. Wasserstein projection estimator
Let $X_{1},\ldots,X_{n}$ be i.i.d. samples from an unknown distribution $\mu^{*}\in\mathcal{P}(\mathbb{R})$. We let $\mu_{n}:=\sum_{i=1}^{n}\frac{1}{n}\delta_{X_{i}}$ be the empirical measure regarded as a random element of $\mathcal{P}(\mathbb{R})$.
@@
Let $p\in(1,\infty)$. Suppose $\mathcal{F}$ is a displacement convex and closed subset of $\mathcal{P}_{p}(\mathbb{R})$ with respect to $\mathcal{W}_{p}$. Given an empirical distribution $\mu_{n}=\sum_{i=1}^{n}\frac{1}{n}\delta_{X_{i}}$, we call
@@
Let $p=2$ and suppose that $\mu^{*}\in\mathcal{P}_{2}(\mathbb{R})$. Let $\mu_{n}=\sum_{i=1}^{n}\frac{1}{n}\delta_{X_{i}}$ where $X_{1},X_{2},\ldots$ are i.i.d. samples from $\mu^{*}$. Then
@@ 3. Monotone density estimation
In this section, we study monotone density estimation under Wasserstein projection. Our main result is Theorem 3.6, which shows that the estimated density is piecewise constant. We begin by formulating an appropriate displacement convex set.
@@
The proof requires a bit of work since we do not impose any regularity condition (other than monotonicity) on the density.
($\Rightarrow$) Suppose $\mu$ has a density $f_{\mu}$ which is non-increasing on $\mathbb{R}_{+}$. Then, $\mu$ is necessarily concentrated on an open interval $I=(0,b)$, where $0<b\leq\infty$, such that $f_{\mu}>0$ on $I$. By [6, Proposition A.18], the quantile function $Q_{\mu}$ is strictly increasing and absolutely continuous on $(0,1)$. Moreover,
@@
$\lim_{\tilde{h}\downarrow 0}\frac{\tilde{h}}{F_{\mu}(x+\tilde{h})-F_{\mu}(x+)}=Q_{\mu}^{\prime}(u)\Rightarrow F_{\mu}^{\prime}(x+)=\frac{1}{Q_{\mu}^{\prime}(F_{\mu}(x)+)},\quad x\in(0,b).$
@@
Let $p>1$, let $\mathcal{F}_{m,p}$ be the set of $\mu\in\mathcal{P}_{p}(\mathbb{R}_{+})$ whose quantile function satisfies the following properties:
@@
Next, we show that $\mathcal{F}_{m,p}$ is closed in $(\mathcal{P}_{p}(\mathbb{R}_{+}),\mathcal{W}_{p})$. By Proposition 2.2 again, it suffices to show that $\mathcal{Q}_{\mathcal{F}_{m,p}}$ is closed in $L^{p}([0,1])$. Suppose $(Q_{n})_{n\geq 1}$ is a sequence in $\mathcal{Q}_{\mathcal{F}_{m,p}}$ converging to some $Q$ in $L^{p}([0,1])$. Since $Q_{n}$ is non-decreasing and convex with $\lim_{u\downarrow 0}Q_{n}(u)=0$, we may extend it to a non-decreasing convex function on $\mathbb{R}$ by letting $Q_{n}(u)=0$ for $u\leq 0$. Since $Q_{n}\rightarrow Q$ in $L^{2}([0,1])$, there exists a subsequence $Q_{n^{\prime}}$ along which $Q_{n^{\prime}}\rightarrow Q$ almost everywhere on $[0,1]$ (and hence on $(-\infty,1)$. By [28, Theorem 10.8], $Q_{n}$ converges pointwise on $\mathbb{R}$ to a convex function $\tilde{Q}$ which is a.e. equal to $Q$ on $(0,1)$. Clearly, $\tilde{Q}$ is non-decreasing. Also, we have $\tilde{Q}(0)=\lim_{n^{\prime}\rightarrow 0}Q_{n^{\prime}}(0)=0$. Furthermore, since $\tilde{Q}$ is continuous at 0 by [28, Theorem 10.1], we have $\tilde{Q}(0)=\lim_{u\rightarrow 0^{+}}\tilde{Q}(u)=0$. Since $Q=\tilde{Q}$ on $(0,\infty)$, we have $Q\in\mathcal{Q}_{\mathcal{F}_{m,p}}$.
@@
$\mathrm{d}\mu(x)=\underbrace{\alpha\mathrm{d}\delta_{0}}_{\text{$\mathrm{d}\mu_{0}(x)$}}+\underbrace{(1-\alpha)I_{(0,1)}\mathrm{d}x}_{\text{$\mathrm{d}\mu_{1}(x)$}},\quad\alpha\in(0,1).$
Note that
@@
Let $p>1$, $\mu\in\mathcal{P}_{p}(\mathbb{R})$ and $\hat{\mu}=\operatorname{proj}_{\mathcal{F}_{m,p}}^{(p)}\mu$.
@@
In particular, if $X_{i}$ are i.i.d. samples from an absolutely continuous distribution $\mu^{*}\in\mathcal{P}_{p}(\mathbb{R})$, the Wasserstein density estimator $\hat{\mu}_{n}=\operatorname{proj}_{\mathcal{F}_{m,p}}\mu_{n}$, where $\mu_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$, is almost surely compactly supported and absolutely continuous.
@@
Case 1. $\nu=\delta_{0}$. Then $Q_{\nu}\equiv 0$. Let $\tilde{\nu}=\mathrm{Unif}(0,\epsilon)\in\mathring{\mathcal{F}}_{m,p}$ which has quantile function $Q_{\tilde{\nu}}(u)=\epsilon u\leq Q_{\mu}(u)$. It follows that
$\begin{split}\mathcal{W}_{p}^{p}(\tilde{\nu},\mu)&=\int_{0}^{1}|\epsilon u-Q_{\mu}(u)|^{p}\mathrm{d}u\\ &<\int_{0}^{1}|0-Q_{\mu}(u)|^{p}\mathrm{d}u\\ &=\mathcal{W}_{p}^{p}(\nu,\mu).\end{split}$
@@
The last equality holds by convexity of $Q_{\nu}$. We have $\tilde{\nu}\in\mathring{\mathcal{F}}_{m,p}$ since $Q_{\nu}$ is strictly increasing. By continuity of $Q_{\nu}$, for $u_{1}$ sufficiently close to $u_{0}$ we have
$Q_{\nu}(u_{1})\frac{u}{u_{1}}\leq\epsilon,\quad u\leq u_{1}.$
For such a choice of $u_{1}$, we have
@@
That is, we cut off the (non-zero) tail of $\nu$ beyond $2M$ and replace that by a mass at 0. By construction, $\tilde{\nu}$ is supported on $[0,2M]$ and is an element of $\mathcal{F}_{m,p}$.
We proceed to show that $\mathcal{W}_{p}(\tilde{\nu},\mu)<\mathcal{W}_{p}(\nu,\mu)$. Let $\pi\in\Pi(\nu,\mu)$ be the monotone coupling between $\nu$ and $\mu$ which is optimal for $\mathcal{W}_{p}(\nu,\mu)$. By the disintegration theorem, we may write
$\pi(\mathrm{d}x,\mathrm{d}y)=\nu(\mathrm{d}x)\pi(\mathrm{d}y|x)$
for some kernel $\pi(\mathrm{d}y|x)$ on $\mathbb{R}_{+}$. Define a coupling $\tilde{\pi}\in\Pi(\tilde{\nu},\mu)$ by
(3.3) $\begin{split}\tilde{\pi}(\mathrm{d}x,\mathrm{d}y)&=I_{[0,2M]}(x)\nu(\mathrm{d}x)\pi(\mathrm{d}y|x)+\tilde{\gamma}(\mathrm{d}x,\mathrm{d}y),\end{split}$
@@
$(1-\nu([0,2M]))\mathrm{d}\delta_{0}(x)$
which is concentrated at 0, and
$\tilde{\mu}(\mathrm{d}y)=\mu(\mathrm{d}y)-\int_{[0,2M]}\nu(\mathrm{d}x)\pi(\mathrm{d}y|x)$
which is supported on $[0,M]$. (Note that both measures have total mass $1-\nu([0,2M])$ by construction.) The coupling $\tilde{\pi}$ is suboptimal for $\mathcal{W}_{p}(\tilde{\mu},\nu)$ but is sufficient for the purposes.
@@
$\begin{split}&\mathcal{W}_{p}^{p}(\tilde{\nu},\mu)\\ &\leq\int_{[0,2M]\times[0,M]}|x-y|^{p}\tilde{\pi}(\mathrm{d}x,\mathrm{d}y)\\ &=\int_{[0,2M]}\left(\int_{[0,M]}|x-y|^{p}\pi(\mathrm{d}y|x)\right)\nu(\mathrm{d}x)+\int_{\{0\}\times[0,M]}|0-y|^{p}\tilde{\gamma}(\mathrm{d}x,\mathrm{d}y)\\ &\leq\int_{[0,2M]}\left(\int_{[0,M]}|x-y|^{p}\pi(\mathrm{d}y|x)\right)\nu(\mathrm{d}x)+M^{p}(1-\nu([0,2M])).\end{split}$
The last inequality holds since
$\int_{\{0\}\times[0,M]}|0-y|^{p}\tilde{\gamma}(\mathrm{d}x,\mathrm{d}y)=\int_{[0,M]}|y|^{p}\tilde{\mu}(\mathrm{d}y)\leq M^{p}\tilde{\nu}([0,M]).$
and $\tilde{\mu}$ has total mass $1-\nu([0,2M])$.
@@
It follows (note $\nu(\{2M\})=0$ and $\nu((2M,\infty))>0$) that
$\int_{[2M,\infty)}\left(\int_{[0,M]}|x-y|^{p}\pi(\mathrm{d}y|x)\right)\nu(\mathrm{d}x)>M^{p}(1-\nu([0,2M])).$
@@
$\mathcal{W}_{p}^{p}(\tilde{\nu},\mu)<\int_{\mathbb{R}_{+}\times[0,M]}|x-y|^{p}\underbrace{\nu(\mathrm{d}x)\pi(\mathrm{d}y|x)}_{\text{$\pi(\mathrm{d}x,\mathrm{d}y)$}}=\mathcal{W}_{p}^{p}(\nu,\mu),$

### 10 Apr 2026 v2

@@ 1. Introduction
Let $X_{1},\ldots,X_{n}$ be independent samples from an unknown distribution $\mu^{*}$. Let $\mathcal{F}$ be a given set of distributions, possibly infinite dimensional, that encodes the shape constraint; it represents our statistical model. The problem is to come up with an estimator $\hat{\mu}_{n}$ of $\mu^{*}$ with values in $\mathcal{F}$. We allow the model $\mathcal{F}$ to be misspecified, so it is possible that $\mu^{*}\notin\mathcal{F}$. Two important cases that will be studied in detail in this paper are (i) the space of non-increasing densities on $\mathbb{R}_{+}$ (the monotone case, see Section 3) and (ii) the space of log-concave densities on $\mathbb{R}$ (the log-concave case, see Section 4).
@@
provided that it is well-posed. In the monotone case, this leads to Grenander’s estimator introduced in [14]. The log-concave case was studied in [41]. Both, and other, shape constraints have been intensively studied in the literature, see for example the monograph [15] and the references [12, 16, 31, 32].
In this paper, we study an alternative approach to the estimation problem, by using optimal transport [33, 39, 40]. For $p\geq 1$, let $\mathcal{W}_{p}(\mu,\nu)$ be the $p$-Wasserstein distance between distributions $\mu$ and $\nu$ on $\mathbb{R}$ (necessary concepts in optimal transport will be reviewed in Section 2.1). Let $\mu_{n}:=\sum_{i=1}^{n}(1/n)\delta_{X_{i}}$ be the empirical distribution of the data. Our Wasserstein projection estimator is defined by
@@
which may be regarded as a special case of the minimum Kantorovich estimator [5]. In order that the Wasserstein projection estimator exists uniquely (see Theorem 2.8), we assume (i) $p>1$, (ii) distributions in $\mathcal{F}$ have finite $p$-th moment, and (iii) $\mathcal{F}$ is closed with respect to $\mathcal{W}_{p}$ and is displacement convex in the sense of McCann [22]. These conditions not only guarantee that (1.2) is a convex optimization problem, but also accommodate a wide variety of shape constraints. In most of the paper, we restrict further to the case $p=2$, where the Wasserstein projection satisfies a desirable Lipschitz property with respect to $\mathcal{W}_{2}$. From this, finite sample performance of the estimator in Wasserstein distance can be established using known results on convergence of empirical distributions (Propositions 2.15 and 4.5).
@@
Although these properties appear to be qualitatively similar to those of the maximum likelihood estimators, there are important differences. For example, the support of the estimated density is generally not equal to the convex hull of the data points, and the set of break points (of intervals where the density is constant or log-affine) may not be a subset of the data points. To give a quick example (see Example 4.4 for details), suppose that the data $\mu_{n}=(1/2)\delta_{-1}+(1/2)\delta_{1}$ is uniformly distributed on $\{-1,1\}$, and $\mathcal{F}$ is the set of log-concave distributions. It is well known that the maximum likelihood estimator returns the uniform distribution $\mathrm{Unif}(-1,1)$. Interestingly, the $2$-Wasserstein projection estimator returns $\mathrm{Unif}(-3/2,3/2)$, which has a wider support. Even in our univariate setting, the proofs of the above structural results are significantly more delicate than those of the classical counterparts. Intuitively, this is because optimal transport incorporates the geometry of the state space. Extending these results to the multivariate setting is interesting but requires different techniques due to the curvature of the Wasserstein space.
@@
Our work contributes to the growing literature of statistical inference based on optimal transport, [^2] This is related to, but different from, inference of transport-based quantities (such as optimal transport maps and Wasserstein barycenters) from samples, and modeling of distributional data. as opposed to classical methods especially maximum likelihood. Recently, an intriguing framework of Wasserstein information geometry has been developed in the parametric setting [20, 3, 2]; of fundamental importance are the Wasserstein information matrix and the resulting Wasserstein–Cramér–Rao inequality [24, 36]. It is interesting to understand to what extent this theory extends to semiparametric and nonparametric settings. Density estimation under Wasserstein distance has been studied by many authors. In particular, minimax estimation of smooth densities in Wasserstein distance was studied in [23, 42], but without shape constraints. Note that smoothness of the density is not assumed in this paper [^3] A non-increasing density may have jump discontinuities, while a log-concave density is continuous (and differentiable except possibly at countably many points) in the interior of its support.. Statistical properties of the Wasserstein projection estimator have been investigated in one-dimensional location-scale models [1] and circular models [25]. Shape constraints were considered in [11], where the author adopted a computational approach and did not prove structural results of the kind obtained here.
The rest of the paper is organized as follows. In Section 2, we introduce the Wasserstein projection estimator after recalling some basic concepts in optimal transport, and establish general properties including consistency and affine equivariance. In Sections 3 and 4, we specialize respectively to monotone and log-concave density estimation, and prove our main results about structural properties of the Wasserstein projection estimator. Section 5 implements tractable discretizations of Wasserstein projection estimation in the monotone and log-concave settings. In Section 6 we conclude and discuss some directions for future research.
@@ 2.1. Wasserstein geometry of univariate distributions
When it is convenient to refer to the boundary values (e.g. when $\mu$ is compactly supported or equivalently $Q_{\mu}$ is bounded), we take $Q_{\mu}(0):=\lim_{u\downarrow 0}Q_{\mu}(u)$ and $Q_{\mu}(1):=\lim_{u\uparrow 1}Q_{\mu}(u)$ which are respectively the $\inf$ and $\sup$ of the support of $\mu$. By construction, we have
@@
where $(Q_{\mu},Q_{\nu})$ denotes the mapping $u\mapsto(Q_{\mu}(u),Q_{\nu}(u))$ from $(0,1)$ to $\mathbb{R}^{2}$. Then $\pi$ is optimal for $\mathcal{W}_{p}(\mu,\nu)$. In fact, by [33, Theorem 2.9], the comonotonic coupling is the unique optimal coupling when $p>1$.
@@
where $\Pi_{t}(x,y):=(1-t)x+ty$ in the linear interpolation between the initial and final locations. By construction, the quantile function of $\mu_{t}$ is $Q_{\mu_{t}}=(1-t)Q_{\mu_{0}}+tQ_{\mu_{1}}$. Thus, a set $\mathcal{F}\subset\mathcal{P}(\mathbb{R})$ is displacement convex in the sense of McCann if and only if $\mathcal{Q}_{\mathcal{F}}$ is convex.
@@
It can be shown that $\mu\in\mathcal{F}$ if and only if $\mu$ has a density $f$ which is bounded below and above on its support: $m\leq f(x)\leq M$. It is easy to see that $\mathcal{F}$ is displacement convex. Note that elements of $\mathcal{F}$ are compactly supported but there is no common support ($\mathcal{F}$ is closed under translation). This family is irregular from the perspective of maximum likelihood but not from that of optimal transport.
@@
Let $\mathcal{F}=\{\delta_{m}:m\in\mathbb{R}\}$ be the set of point masses. Clearly, $\mathcal{F}$ is displacement convex. Let $\mu=(1/2)\delta_{0}+(1/2)\delta_{1}$. It is easy to verify that
@@
$\begin{split}\mathcal{W}_{p}^{p}(T_{\#}\nu,T_{\#}\mu)=\int_{0}^{1}|Q_{T_{\#}\nu}(u)-Q_{T_{\#}\mu}(u)|^{p}\mathrm{d}u=|b|^{p}\mathcal{W}_{p}^{p}(\nu,\mu).\end{split}$
@@ 2.2. Wasserstein projection estimator
Let $X_{1},\ldots,X_{n}$ be i.i.d. samples from an unknown distribution $\mu^{*}\in\mathcal{P}(\mathbb{R})$. We let $\mu_{n}:=\sum_{i=1}^{n}(1/n)\delta_{X_{i}}$ be the empirical measure regarded as a random element of $\mathcal{P}(\mathbb{R})$.
@@
Let $p\in(1,\infty)$. Suppose $\mathcal{F}$ is a displacement convex and closed subset of $\mathcal{P}_{p}(\mathbb{R})$ with respect to $\mathcal{W}_{p}$. Given an empirical distribution $\mu_{n}=\sum_{i=1}^{n}(1/n)\delta_{X_{i}}$, we call
@@
Recall that we write $\operatorname{proj}_{\mathcal{F}}^{(2)}=\operatorname{proj}_{\mathcal{F}}$.
@@
Let $p=2$ and suppose that $\mu^{*}\in\mathcal{P}_{2}(\mathbb{R})$. Let $\mu_{n}=\sum_{i=1}^{n}(1/n)\delta_{X_{i}}$ where $X_{1},X_{2},\ldots$ are i.i.d. samples from $\mu^{*}$. Then
@@ 3. Monotone density estimation
In this section, we study monotone density estimation under Wasserstein projection. Our main result is Theorem 3.6, which shows that the estimated density is piecewise constant. We begin by formulating an appropriate displacement convex set for the monotonicity constraint.
@@
The proof requires a bit of work since we do not impose any regularity condition, other than monotonicity, on the density.
($\Rightarrow$) Suppose $\mu$ has a density $f_{\mu}$ which is non-increasing on $\mathbb{R}_{+}$. Then, $\mu$ is necessarily concentrated on an open interval $I=(0,b)$, where $0<b\leq\infty$, such that $f_{\mu}>0$ on $I$. By [6, Proposition A.18], the quantile function $Q_{\mu}$ of $\mu$ is strictly increasing and absolutely continuous on $(0,1)$. Moreover, the chordal slope of $Q_{\mu}$ is given by
@@
$\lim_{\tilde{h}\downarrow 0}\frac{\tilde{h}}{F_{\mu}(x+\tilde{h})-F_{\mu}(x+)}=Q_{\mu}^{\prime}(u+)\Rightarrow F_{\mu}^{\prime}(x+)=\frac{1}{Q_{\mu}^{\prime}(F_{\mu}(x)+)},\quad x\in(0,b).$
@@
For $p>1$, let $\mathcal{F}_{m,p}$ be the set of $\mu\in\mathcal{P}_{p}(\mathbb{R}_{+})$ whose quantile function satisfies the following properties:
@@
Next, we show that $\mathcal{F}_{m,p}$ is closed in $(\mathcal{P}_{p}(\mathbb{R}_{+}),\mathcal{W}_{p})$. By Proposition 2.2 again, it suffices to show that $\mathcal{Q}_{\mathcal{F}_{m,p}}$ is closed in $L^{p}([0,1])$. Suppose $(Q_{n})_{n\geq 1}$ is a sequence in $\mathcal{Q}_{\mathcal{F}_{m,p}}$ converging to some $Q$ in $L^{p}([0,1])$. Since $Q_{n}$ is non-decreasing and convex with $\lim_{u\downarrow 0}Q_{n}(u)=0$, we may extend it to a non-decreasing convex function on $\mathbb{R}$ by letting $Q_{n}(u)=0$ for $u\leq 0$. Since $Q_{n}\rightarrow Q$ in $L^{p}([0,1])$, there exists a subsequence $Q_{n^{\prime}}$ along which $Q_{n^{\prime}}\rightarrow Q$ almost everywhere on $[0,1]$ (and hence on $(-\infty,1)$. By [29, Theorem 10.8], $Q_{n}$ converges pointwise on $\mathbb{R}$ to a convex function $\tilde{Q}$ which is equal to $Q$ on $(0,1)$. Clearly, $\tilde{Q}$ is non-decreasing. Also, we have $\tilde{Q}(0)=\lim_{n^{\prime}\rightarrow\infty}Q_{n^{\prime}}(0)=0$. Furthermore, since $\tilde{Q}$ is continuous at 0 by [29, Theorem 10.1], we have $\tilde{Q}(0)=\lim_{u\rightarrow 0^{+}}\tilde{Q}(u)$. Since $Q=\tilde{Q}$ on $(0,\infty)$, we have $Q\in\mathcal{Q}_{\mathcal{F}_{m,p}}$.
@@
$\mathrm{d}\mu(x)=\underbrace{\alpha\mathrm{d}\delta_{0}}_{\text{$\mathrm{d}\mu_{0}(x)$}}+\underbrace{(1-\alpha)I_{(0,1)}\mathrm{d}x}_{\text{$\mathrm{d}\mu_{1}(x)$}},$
where $\alpha\in(0,1)$. Note that
@@
Let $p>1$, $\mu\in\mathcal{P}_{p}(\mathbb{R}_{+})$ and $\hat{\mu}=\operatorname{proj}_{\mathcal{F}_{m,p}}^{(p)}\mu$.
@@
In particular, if $X_{i}$ are i.i.d. samples from an absolutely continuous distribution $\mu^{*}\in\mathcal{P}_{p}(\mathbb{R})$, the Wasserstein density estimator $\hat{\mu}_{n}=\operatorname{proj}_{\mathcal{F}_{m,p}}\mu_{n}$, where $\mu_{n}=\sum_{i=1}^{n}(1/n)\delta_{X_{i}}$, is almost surely compactly supported and absolutely continuous.
@@
Case 1. $\nu=\delta_{0}$. Then $Q_{\nu}\equiv 0$. Let $\tilde{\nu}=\mathrm{Unif}(0,\epsilon)\in\mathring{\mathcal{F}}_{m,p}$ which has quantile function $Q_{\tilde{\nu}}(u)=\epsilon u$. Since $0=Q_{\nu}<Q_{\tilde{\nu}}\leq Q_{\mu}$ on $(0,1)$, we have
$\begin{split}\mathcal{W}_{p}^{p}(\tilde{\nu},\mu)=\int_{0}^{1}|\epsilon u-Q_{\mu}(u)|^{p}\mathrm{d}u<\int_{0}^{1}|0-Q_{\mu}(u)|^{p}\mathrm{d}u=\mathcal{W}_{p}^{p}(\nu,\mu).\end{split}$
@@
The last equality holds by convexity of $Q_{\nu}$. It is easy to verify that $\tilde{\nu}\in\mathring{\mathcal{F}}_{m,p}$. By continuity of $Q_{\nu}$, for $u_{1}$ sufficiently close to $u_{0}$ we have $Q_{\nu}(u_{1})\leq\epsilon$. For such a choice of $u_{1}$, we have
@@
That is, we cut off the (non-zero) tail of $\nu$ beyond $2M$ and replace that by a mass at 0.666Note that here it is possible that $\mu$ charges the point 0. If not, we may use the construction in (i) to remove the mass at 0. By construction, $\tilde{\nu}$ is supported on $[0,2M]$ and is an element of $\mathcal{F}_{m,p}$.
We proceed to show that $\mathcal{W}_{p}(\tilde{\nu},\mu)<\mathcal{W}_{p}(\nu,\mu)$. Let $\pi\in\Pi(\nu,\mu)$ be the comonotonic coupling between $\nu$ and $\mu$ which is optimal for $\mathcal{W}_{p}(\nu,\mu)$. By the disintegration theorem, we may write
$\pi(\mathrm{d}x,\mathrm{d}y)=\nu(\mathrm{d}x)\pi(\mathrm{d}y\mid x)$
for some kernel $\pi(\mathrm{d}y\mid x)$ on $\mathbb{R}_{+}$. Define a coupling $\tilde{\pi}\in\Pi(\tilde{\nu},\mu)$ by
(3.4) $\begin{split}\tilde{\pi}(\mathrm{d}x,\mathrm{d}y)&=I_{[0,2M]}(x)\nu(\mathrm{d}x)\pi(\mathrm{d}y\mid x)+\tilde{\gamma}(\mathrm{d}x,\mathrm{d}y),\end{split}$
@@
$(1-\nu([0,2M]))\mathrm{d}\delta_{0}(x),$
which is concentrated at 0, and the remainder
$\tilde{\mu}(\mathrm{d}y)=\mu(\mathrm{d}y)-\int_{[0,2M]}\nu(\mathrm{d}x)\pi(\mathrm{d}y\mid x),$
which has the same mass $1-\nu([0,2M])$ and is supported on $[0,M]$. The coupling $\tilde{\pi}$ is suboptimal for $\mathcal{W}_{p}(\tilde{\nu},\mu)$ but is sufficient for the purposes.
@@
$\begin{split}&\mathcal{W}_{p}^{p}(\tilde{\nu},\mu)\\ &\leq\int_{[0,2M]\times[0,M]}|x-y|^{p}\tilde{\pi}(\mathrm{d}x,\mathrm{d}y)\\ &=\int_{[0,2M]}\left(\int_{[0,M]}|x-y|^{p}\pi(\mathrm{d}y\mid x)\right)\nu(\mathrm{d}x)+\int_{\{0\}\times[0,M]}|0-y|^{p}\tilde{\gamma}(\mathrm{d}x,\mathrm{d}y)\\ &\leq\int_{[0,2M]}\left(\int_{[0,M]}|x-y|^{p}\pi(\mathrm{d}y\mid x)\right)\nu(\mathrm{d}x)+M^{p}(1-\nu([0,2M])).\end{split}$
@@
It follows that
$\int_{[2M,\infty)}\left(\int_{[0,M]}|x-y|^{p}\pi(\mathrm{d}y\mid x)\right)\nu(\mathrm{d}x)>M^{p}(1-\nu([0,2M])).$
@@
$\mathcal{W}_{p}^{p}(\tilde{\nu},\mu)<\int_{\mathbb{R}_{+}\times[0,M]}|x-y|^{p}\underbrace{\nu(\mathrm{d}x)\pi(\mathrm{d}y\mid x)}_{\text{$\pi(\mathrm{d}x,\mathrm{d}y)$}}=\mathcal{W}_{p}^{p}(\nu,\mu),$